# Vibrational Fingerprints of Strained Polymers: A Spectroscopic Pathway to Mechanical State Prediction

### v1

@@ Introduction
In this paper, we investigate the vibrational properties of crosslinked epoxy networks under uniaxial tension using MD simulations and experimental in-situ tensile tests on film samples made from bisphenol A diglycidyl ether (DGEBA) and an aliphatic hardener to verify the simulation results. We analyze redshifts in the para-phenylene stretching modes and bending vibrations of aromatic hydrogens under load. For the MD simulations, two epoxy systems, each composed of a different resin monomer and either an aliphatic or an aromatic hardener, are examined with both the classical OPLS-AA force field and the ML potential MACE-OFF23. In addition, we train a machine-learned dipole model on chemically representative epoxy fragments and apply it to fragment-level simulations to provide IR-intensity information complementary to velocity-based spectra. Together, these approaches bridge the gap between molecular-scale structure and vibrational response under strain in crosslinked polymer networks. By combining charge-weighted VACF analysis, ML potentials, and SAGPR-derived dipoles, we resolve strain-induced modifications to aromatic vibrational signatures. This framework reveals the molecular origins of spectral changes under stress, with implications for diagnostics, failure prediction, and the design of next-generation polymer composites.
@@
We used LAMMPS [43] to model the crosslinking reactions and bulk polymer formation of the BFDGE-DETDA and DGEBA-DETA epoxy networks. Each simulation box contained 160 resin molecules (BFDGE or DGEBA) and 80 DETDA or 64 DETA hardener molecules, yielding total system sizes of 9360 (BFDGE-DETDA) and 9120 (DGEBA-DETA) atoms. We equilibrated the systems in the NPT ensemble for 2 $\mathrm{n}\mathrm{s}$, followed by a crosslinking procedure performed over 2 $\mathrm{n}\mathrm{s}$ at 600 K using the LAMMPS REACTION package [44]. We used a time step of 0.5 $\mathrm{f}\mathrm{s}$, and applied the NVE/limit integrator for 500 $\mathrm{f}\mathrm{s}$ to constrain reactive-site displacements to 0.0015 Å per $\mathrm{f}\mathrm{s}$.
We applied Nosé–Hoover thermostat and barostat coupling with damping constants of 100 $\mathrm{f}\mathrm{s}$ and 1000 $\mathrm{f}\mathrm{s}$, respectively, at a pressure of 1 $\mathrm{a}\mathrm{t}\mathrm{m}$ [45]. Periodic boundary conditions were imposed in all directions. We computed Lennard-Jones interactions with an 8 Å cutoff and short-range electrostatics with a 12 Å cutoff. Long-range electrostatics were treated using the particle–particle particle–mesh method with a $k$-space accuracy of $1.0\times 10^{-4}$ [46].
@@
We generated conformational ensembles of each fragment using replica exchange molecular dynamics (REMD) with the OPLS-AA force field. Temperature windows spanned 300–580 K in 40 K increments, ensuring coverage of the relevant configurational space. Each REMD trajectory ran for 500 ns, providing sufficient sampling of both local torsional motions and large-scale rearrangements. We extracted structures for QM calculations exclusively from the 300 K replica to retain thermally accessible conformers at ambient conditions. To reduce redundancy, we applied farthest-point clustering in the space of 17 dihedral descriptors, which reflect the dominant internal degrees of freedom of the fragments. This yielded 5000 unique structures per fragment, producing a diverse and balanced dataset.
@@
Table 1: Vibrational wavenumbers derived from the charge-weighted velocity autocorrelation function (VACFq) for the OPLS-AA and MACE-OFF23 force fields at different strain levels ($\epsilon$) in the BFDGE-DETDA system.
@@
By isolating the aromatic contributions (Figure 3), we found that the resulting spectra selectively emphasize aromatic vibrational modes. The analysis focused on characteristic features: the $\mathrm{C}_{1}\text{--}\mathrm{C}_{4}$ stretching modes (shown in red), comprising a symmetric stretch near 1650 cm-1 and an asymmetric stretch around 1550 cm-1, and the $\mathrm{C}_{\mathrm{ar}}\text{--}\mathrm{H}$ bending modes (shown in blue), including an out-of-plane (oop) vibration at approximately 850 cm-1 and an in-plane (ip) vibration near 990 cm-1 [58]. This decomposition sharpened peak assignment and clarified the molecular origins of the aromatic features within the epoxy network.
@@
The mechanical properties, including ultimate tensile strength and strain at break, showed no systematic dependence on strain rate. In contrast, the $\nu_{\mathrm{C_{1}-C_{4}}}^{\mathrm{sym}}$ and $\nu_{\mathrm{C_{1}-C_{4}}}^{\mathrm{asym}}$ stretching vibrations exhibited a tendency toward larger redshift slopes at slower strain rates, whereas the $\nu_{\mathrm{C-H}}^{\mathrm{oop}}$ bending vibration showed no reproducible trend. The variability across experiments attributed to sample-to-sample differences arising from the manual preparation of thin films, such as slight variations in thickness or the presence of micro-defects. The data shown in Figure 5 represent a typical test at $\dot{\epsilon}=0.5$ mm/min, while additional results at different strain rates and the corresponding strain rate dependency are provided in Supplementary Section S3.
@@
Stronger chain alignment was observed in the DGEBA–DETA system than in BFDGE–DETDA, as reflected by the steeper slope of $d_{\mathrm{chain}}$ and $d_{\mathrm{F}}$ in response to strain. This structural adaptation induced elongation of the para-phenylene group, quantified by the distance between $\mathrm{C}_{1}$ and $\mathrm{C}_{4}$ atoms, $d_{\mathrm{C}_{1}\text{--}\mathrm{C}_{4}}$, which correlates with the redshift of its vibrational modes in accordance with Badger’s rule [59]. The linear relationship observed in Figure 6 b) confirms that strain-induced molecular elongation of the aromatic ring directly contributes to the frequency shifts in the vibrational spectra. This trend highlights the increased stretching of para-phenylene units in the DGEBA–DETA system, consistent with its stronger overall alignment behavior.
Although both epoxy systems exhibited broadly similar trends, subtle but meaningful differences emerged. In DGEBA–DETA, chain alignment was approximately 20 % more pronounced than in BFDGE–DETDA, as indicated by the slope of $d_{\mathrm{F}}$. This stronger alignment was accompanied by in the increase of $d_{\mathrm{C}_{1}\text{--}\mathrm{C}_{4}}$. However, the corresponding redshift of the para-phenylene stretching modes did not scale proportionally. This discrepancy suggests that the spectral response is governed not only by geometric elongation but also by local network flexibility, anharmonic coupling, and vibrational mode mixing.
@@
To assign specific peaks within the spectra, we applied targeted ACF–FFT workflows. For the symmetric stretching vibration of the para-phenylene group, we sampled the $\mathrm{C}_{1}$$\mathrm{C}_{4}$ distance and computed its power spectrum (red). Analogously, we isolated the out-of-plane bending motion of the aromatic hydrogens by calculating fluctuations of the corresponding distance coordinate and transforming them into the frequency domain (blue). This approach allowed us to unambiguously assign peaks to distinct vibrational modes, providing a direct structural interpretation of the spectral features.
@@ Acknowledgments
The supplementary information includes: Supplementary Text, Figure S1 and S2, Tables S1 and S2.

### v2

@@ Introduction
In this paper, we investigate the vibrational properties of crosslinked epoxy networks under uniaxial tension using MD simulations and experimental in-situ tensile tests on film samples made from bisphenol A diglycidyl ether (DGEBA) and an aliphatic hardener to verify the simulation results. We analyze redshifts in the para-phenylene stretching modes and bending vibrations of aromatic hydrogens under load [40]. For the MD simulations, two epoxy systems, each composed of a different resin monomer and either an aliphatic or an aromatic hardener, are examined with both the classical OPLS-AA force field and the ML potential MACE-OFF23. In addition, we train a machine-learned dipole model on chemically representative epoxy fragments and apply it to fragment-level simulations to provide IR-intensity information complementary to velocity-based spectra. Together, these approaches bridge the gap between molecular-scale structure and vibrational response under strain in crosslinked polymer networks. By combining charge-weighted VACF analysis, ML potentials, and SAGPR-derived dipoles, we resolve strain-induced modifications to aromatic vibrational signatures. This framework reveals the molecular origins of spectral changes under stress, establishing a foundation for non-destructive diagnostics and reliability assessment in polymer networks.
@@
We used LAMMPS [44] to model the crosslinking reactions and bulk polymer formation of the BFDGE-DETDA and DGEBA-DETA epoxy networks. Each simulation box contained 160 resin molecules (BFDGE or DGEBA) and 80 DETDA or 64 DETA hardener molecules, yielding total system sizes of 9360 (BFDGE-DETDA) and 9120 (DGEBA-DETA) atoms. We equilibrated the systems in the NPT ensemble for 2 $\mathrm{n}\mathrm{s}$, followed by a crosslinking procedure performed over 2 $\mathrm{n}\mathrm{s}$ at 600 K using the LAMMPS REACTION package [45]. In case of a reaction happening, we used a time step of 0.5 $\mathrm{f}\mathrm{s}$, and applied the NVE/limit integrator for 500 $\mathrm{f}\mathrm{s}$ to constrain reactive-site displacements to 0.0015 Å per $\mathrm{f}\mathrm{s}$.
We applied Nosé–Hoover thermostat and barostat coupling with damping constants of 100 $\mathrm{f}\mathrm{s}$ and 1000 $\mathrm{f}\mathrm{s}$, respectively, at a pressure of 1 $\mathrm{a}\mathrm{t}\mathrm{m}$ [46]. Periodic boundary conditions were imposed in all directions. We computed Lennard-Jones interactions with an 8 Å cutoff and short-range electrostatics with a 12 Å cutoff. Long-range electrostatics were treated using the particle–particle particle–mesh method with a (relative) $k$-space accuracy of $1.0\times 10^{-4}$ [47].
@@
We generated conformational ensembles of each fragment using replica exchange molecular dynamics (REMD) with the OPLS-AA force field. Temperature windows spanned 300–580 K in 40 K increments, ensuring coverage of the relevant configurational space. Each REMD trajectory ran for 500 ns, providing sufficient sampling of both local torsional motions and large-scale rearrangements. We extracted structures for QM calculations exclusively from the 300 K replica to retain thermally accessible conformers at ambient conditions. To reduce redundancy, we applied farthest-point sampling in the space of 17 dihedral descriptors, which reflect the dominant internal degrees of freedom of the fragments. This yielded 5000 unique structures per fragment, producing a diverse and balanced dataset.
@@
Table 1: Vibrational wavenumbers derived from the charge-weighted velocity autocorrelation function (VACFq) for the OPLS-AA and MACE-OFF23 force fields at different strain levels ($\epsilon$) in the BFDGE-DETDA system. The indices (1–7) correspond to the peak positions identified in Figure 2b.
@@
By isolating the aromatic contributions through evaluation of the charge-weighted velocity autocorrelation function using only the velocities of the aromatic atoms (Figure 3), we obtained spectra that selectively emphasize aromatic vibrational modes. The analysis focused on characteristic features: the $\mathrm{C}_{1}\text{--}\mathrm{C}_{4}$ stretching modes (shown in red), comprising a symmetric stretch near 1650 cm-1 and an asymmetric stretch around 1550 cm-1, and the $\mathrm{C}_{\mathrm{ar}}\text{--}\mathrm{H}$ bending modes (shown in blue), including an out-of-plane (oop) vibration at approximately 850 cm-1 and an in-plane (ip) vibration near 990 cm-1 [59]. This decomposition sharpened peak assignment and clarified the molecular origins of the aromatic features within the epoxy network.
@@
The mechanical properties, including ultimate tensile strength and elongation at break, showed no systematic dependence on strain rate. In contrast, the $\nu_{\mathrm{C_{1}-C_{4}}}^{\mathrm{sym}}$ and $\nu_{\mathrm{C_{1}-C_{4}}}^{\mathrm{asym}}$ stretching vibrations exhibited a tendency toward larger redshift slopes at slower strain rates, whereas the $\nu_{\mathrm{C-H}}^{\mathrm{oop}}$ bending vibration showed no reproducible trend. The variability across experiments attributed to sample-to-sample differences arising from the manual preparation of thin films, such as slight variations in thickness or the presence of micro-defects. The data shown in Figure 5 represent a typical test at $\dot{\epsilon}=0.5$ mm/min, while additional results at different strain rates and the corresponding strain rate dependency are provided in Supplementary Section S3.
@@
Stronger chain alignment was observed in the DGEBA–DETA system than in BFDGEDETDA, as reflected by the steeper slope of $d_{\mathrm{chain}}$ and $d_{\mathrm{F}}$ in response to strain. This structural adaptation induced elongation of the para-phenylene group, quantified by the distance between $\mathrm{C}_{1}$ and $\mathrm{C}_{4}$ atoms, $d_{\mathrm{C}_{1}\text{--}\mathrm{C}_{4}}$, which correlates with the redshift of its vibrational modes in accordance with Badger’s rule [60]. The linear relationship observed in Figure 6 b) confirms that strain-induced molecular elongation of the aromatic ring directly contributes to the frequency shifts in the vibrational spectra. This trend highlights the increased stretching of para-phenylene units in the DGEBA–DETA system, consistent with its stronger overall alignment behavior.
Although both epoxy systems exhibited broadly similar trends, subtle but meaningful differences emerged. In DGEBA–DETA, chain alignment was approximately 20 % more pronounced than in BFDGE–DETDA, as indicated by the slope of $d_{\mathrm{F}}$. This stronger alignment was accompanied by an increase of $d_{\mathrm{C}_{1}\text{--}\mathrm{C}_{4}}$. However, the corresponding redshift of the para-phenylene stretching modes did not scale proportionally. This discrepancy suggests that the spectral response is governed not only by geometric elongation but also by local network flexibility, anharmonic coupling, and vibrational mode mixing.
@@
To assign specific peaks within the spectra, we applied targeted autocorrelation function–fast Fourier transform (ACF–FFT). For the symmetric stretching vibration of the para-phenylene group, we sampled the $\mathrm{C}_{1}$$\mathrm{C}_{4}$ distance and computed its power spectrum (red curve in Fig. 9). Analogously, we isolated the out-of-plane bending motion of the aromatic hydrogens by calculating fluctuations of the corresponding distance coordinate and transforming them into the frequency domain (blue curve in Fig. 9). This approach allowed us to unambiguously assign peaks to distinct vibrational modes, providing a direct structural interpretation of the spectral features.
@@ Acknowledgments
The supplementary information includes: Supplementary Text, Figures S1–S5, Tables S1 and S2.